6) A graphic design studio is considering three new computers. The first model, A, costs $5000 Model B and C cost $3000 and $1000 respectively. If each customer provides $50 of revenue and variable costs are $20/customer, find the number of customers required for each model to break even. (5 points)
Answer #1

As per simple break even formula

Q= F/(P-V) where F= Toatal fixed cost, P=sailing prise / unit , V = variable cost/ unit

In given case Fixed cost in $ Revenue in $ Variable cost $ Q=F/(P-V) BEP units

A 5000.00 50.00 20.00 5000/(50-20) 166.66 approx 167

B 3000.00 50.00 20.00 3000/(50-20) 100.00

C 1000.00 50.00 20.00 1000(50-20) 33.33 approx 33

---------------------------------------------------------------------------------------------------------------------------------------------------------------- Verifying result Total inflow (QxP) = total outflow (Fixd cost+ variable cost)

case A 167x 50.00 = 8350.00 5000.00+20.00X167=8340.00

case B 100X50.00= 5000.00 3000.00+20.00x100=5000.00

case C 33 x50.00= 1650.00 1000.00+20.00x33= 1660.00

Difference error of 10$ exist due to desimal correction in case A & C and result are verified.

--------------------------------------------------------------------------------------------------------------------------------------------------------------------